# FoundPose: Unseen Object Pose Estimation with Foundation Features

### v1

@@ 3.2 Template-based object representation
Template generation. Given a texture-mapped 3D object model, we render $n$ RGB-D templates showing the model under different orientations. The orientations are sampled to uniformly cover the $SO(3)$ group of 3D rotations [1], and the model is rendered using a standard rasterization technique [81] with a black background and fixed lighting. The size of templates is $S\times S$ pixels and the objects are rendered such as the longer side of their 2D bounding box is $\delta S$ pixels long, with $\delta<1$. At inference, we generate crops of the query image with the same size and padding (to allow for errors of segmentation masks around which we crop the image).
@@
Patch descriptors registered in 3D. For each RGB-D template with an index $t\in\{1,\dots,n\}$, we split the RGB channels into $m$ non-overlapping patches of $14\times 14$ pixels and calculate their patch descriptors $\{\mathbf{p}_{t,i}\}_{i=1}^{m}$. A patch descriptor is calculated as $\mathbf{p}_{t,i}=\phi_{d}(\mathbf{p}^{\prime}_{t,i})$, where $\mathbf{p}^{\prime}_{t,i}$ is the raw patch descriptor extracted by DINOv2 and $\phi_{d}:\mathbb{R}^{r}\mapsto\mathbb{R}^{d}$ projects the $r$ dimensional raw descriptor to the top $d$ PCA components calculated from valid patch descriptors of all $n$ templates. A patch is considered valid if its 2D center falls inside the object mask and the PCA-based dimensionality reduction is applied to increase efficiency. Then we represent a template $t$ by a set $T_{t}=\{(\mathbf{p}_{t,j},\mathbf{x}_{j})\;|\;j\in M\}$, where $M$ are indices of valid patches, $\mathbf{x}_{j}$ is a 3D location in the object model space whose 2D projection is at the center of patch $j$. The 3D locations, which are calculated from the depth channel of the template and known camera intrinsics, enable establishing 2D-3D correspondences at inference.
@@
Bag-of-words descriptors. At onboarding, we also pre-calculate bag-of-words descriptors of all templates to enable efficient template retrieval at inference using the classical bag-of-words image retrieval technique [84, 72], which mimics text-retrieval systems with the analogy of visual words. Specifically, we define visual words as the centroids of $k$-means clusters of patch descriptors extracted from all templates of an object. To calculate the bag-of-words descriptor of a template $t$, we assign patch descriptors from the template representation $T_{t}$ to the nearest visual words and describe the template by a vector $\mathbf{b}_{t}=(b_{1},b_{2},\dots,b_{k})$. This vector consists of weighted word frequencies defined as $b_{i}=\frac{n_{i,t}}{n_{t}}\log\frac{N}{n_{i}}$, where $n_{i,t}$ is the number of occurrences of word $i$ in template $t$, $n_{t}$ is the total number of words in template $t$, $n_{i}$ is the number of occurrences of word $i$ in all $N$ templates. The first term ($\frac{n_{i,t}}{n_{t}}$) weights words that occur often in a particular template and therefore describe the template well, while the second term ($\log\frac{N}{n_{i}}$) downweights words that occur often in any template. As visual words generated by clustering may suffer from quantization errors, we follow [72, 94] and soft-assign each patch descriptor to several nearest words with weights defined by $\exp-\frac{d^{2}}{2\sigma^{2}}$, where $d$ is the Euclidean distance of the descriptor from the word and $\sigma$ is a parameter of the method.
@@ 3.3 Template retrieval by bag-of-words matching
Perspective cropping. At inference, we start by cropping the image region around a given object segmentation mask. To minimize perspective distortion and achieve a crop that resembles a template, we generate the crop by warping the query image to a virtual pinhole camera focused at the segmentation mask. The virtual camera is constructed such as its optical axis is passing through the center of the 2D bounding box of the mask, the viewport size is $S\times S$ pixels, and the longer side of the warped 2D bounding box is $\delta S$ pixels long.
Retrieving similar templates. To retrieve a small set of templates, we calculate the bag-of-words descriptor of the crop (as in Sec. 3.2) and calculate its cosine similarity (i.e., normalized scalar product) with bag-of-words descriptors of all object templates. We select $h$ templates with the highest cosine similarity, which provide approximate hypotheses on the object orientation for the subsequent pose estimation stage. This retrieval technique is robust to partial occlusions and remarkably efficient.
Since the bag-of-words descriptor represents an image as a bag of unordered visual words, the retrieval can be typically improved by re-ranking the results with a spatial verification stage [71]. However, in our case, a similar verification is implicitly done by the subsequent PnP-RANSAC (spatially consistent correspondences are expected to yield a better pose estimate) and bags of words are not completely unordered as the used patch descriptors from an intermediate DINOv2 layer, from which the words are constructed, contains 2D positional information.
Figure 3: Visualization of DINOv2 patch descriptors. Shown are top three PCA components of patch descriptors extracted at different layers of DINOv2 ViT-L[14] from templates of a textured object from YCB-V[100] and a symmetric and texture-less object from T-LESS [31]. As observed in[2] and also clearly visible in these visuals, patch descriptors contain gradually less positional and more semantic information when going from shallower to deeper layers. FoundPose performs the best with descriptors from layer 18, which presumably provides the right information mix. The positional information is important for producing geometrically consistent correspondences when the semantic information is not discriminative due to symmetries or a lack of texture.
@@ 3.4 Pose estimation from 2D-3D correspondences
Crop-to-template patch matching. For each retrieved template $t$, we match patch descriptors from the image crop to the nearest (in terms of the Euclidean distance) descriptors from $T_{t}$ and establish 2D-3D correspondences $C_{t}=\left\{\left(\mathbf{u}_{i},\mathbf{x}_{i}\right)\right\}_{i=1}^{m}$, where $\mathbf{u}_{i}$ is the 2D center of a patch in the query image and $\mathbf{x}_{i}$ is the 3D location associated with the matched patch descriptor from $T_{t}$.
Establishing 2D-3D correspondences by crop-to-template patch matching is a considerably simpler problem than exhaustive matching against patch descriptors from all templates, which would be necessary without the template retrieval. Moreover, we demonstrate that the template-based approach can effectively handle arbitrary objects, including challenging objects with symmetries and without a significant texture. The ambiguity of 2D-3D correspondences [30], for which such objects are notoriously known, is eliminated by (1) restricting the candidate descriptors to ones from a single template and (2) using patch descriptors from intermediate layers of DINOv2 that contain both semantic and 2D positional information [2] (Fig. 3). The positional information is the key for producing geometrically consistent correspondences when the semantic information is not discriminative due to symmetries or a lack of texture.
Pose fitting. An object pose $(\mathbf{R}_{t},\mathbf{t}_{t})$, defined by a 3D rotation $\mathbf{R}_{t}$ and 3D translation $\mathbf{t}_{t}$ from the model space to the camera space, is estimated for each retrieved template $t$ from 2D-3D correspondences $C_{t}$ by solving the Perspective-n-Point (PnP) problem. We solve this problem by the EPnP algorithm [43] combined with the RANSAC fitting scheme [20] for robustness. In this scheme, PnP is solved repeatedly on a randomly sampled minimal set of 4 correspondences and the final output is defined by the pose hypothesis with the highest quality, which we define by the number of inlier correspondences [20]. From the set of $h$ poses estimated from the $h$ retrieved templates, the pose with the highest quality is selected as the final coarse pose estimate $(\mathbf{R}_{c},\mathbf{t}_{c})$.
@@ 3.5 Featuremetric pose refinement
Featuremetric alignment. The coarse pose $(\mathbf{R}_{c},\mathbf{t}_{c})$ is further refined by Levenberg-Marquardt (L-M) [44, 54], an iterative non-linear optimization algorithm. Each L-M iteration updates the pose parameters by a gradient-descent step minimizing a real-valued cost, while adaptively selecting between the first and second order gradients depending on the cost value. We initialize the optimization with the coarse pose $(\mathbf{R}_{c},\mathbf{t}_{c})$ and obtain a refined pose $(\mathbf{R}_{r},\mathbf{t}_{r})$ by minimizing a featuremetric error:
@@
where $\rho$ is a robust cost function by Barron [5], $(\mathbf{p}_{i},\mathbf{x}_{i})\in T_{t}$ is a descriptor and the corresponding 3D location of patch $i$ from template $t$, and $\pi:\mathbb{R}^{3}\mapsto\mathbb{R}^{2}$ represents the 2D projection. The feature map $\textbf{F}_{q}\in\mathbb{R}^{a\times a\times d}$ is obtained by stacking the patch descriptors of the query image and is sampled with bilinear interpolation at normalized 2D projections. The spatial resolution of the feature map is $a\times a$ with $a=S/s$, where $s\times s$ is the patch size. Note that a similar featuremetric alignment was applied in several works [95, 79], typically in combination with features trained specifically for L-M. In our work, we apply optimization directly on DINOv2 features, without any training.
Reducing 2D-3D discrepancy. The 2D-3D correspondences are established by matching patch descriptors extracted on a regular grid in the query image given by the 2D patch centers. Such a coarse sampling may reduce the accuracy of pose estimates as the centers of matched patches may not be aligned with 2D projections of the same 3D points. To compensate for the potential discrepancy, we refine the pose estimates by aligning template patches to their optimal locations in the query image. Another possible approach to reduce the discrepancy is to increase spatial density of the extracted patch descriptors, which could be achieved, e.g., by splitting the image into overlapping patches [2]. However, this approach reduces efficiency while not improving accuracy of our method.
Figure 4: Example FoundPose results on HB, LM-O, IC-BIN, TUD-L, ITODD and T-LESS datasets. Each example shows the query image crop with the CNOS mask in white (top left), retrieved templates (middle row), matched patch descriptors of the crop and the template that led to a pose with the highest number of inlier correspondences (bottom row), and the ground-truth pose contour in red, the coarse pose in blue and the refined pose in green (top right). These examples demonstrate that FoundPose can handle a broad range or objects, including textured, texture-less and symmetric ones.
@@ 4.1 Experimental setup
Method Pose refinement
@@
01 FoundPose 39.6 33.8 46.7 23.9 20.4 50.8 45.2 37.2
01.7
2 GenFlow [57]
25.0
21.5 30.0 16.8 15.4
28.3
27.7
23.5
03.8
3 MegaPose [41]
22.9
17.7
25.8
15.2
10.8
25.1
28.1
20.8
15.5
4 OSOP [82] 31.2
49.2 33.2
5 ZS6D [3]
29.8
21.0
32.4
@@ 5 ZS6D [3]
06 FoundPose Featuremetric
39.5
39.6
56.7
28.3
26.2
58.5
49.7
42.6
02.6
07 FoundPose MegaPose 55.4 51.0 63.3 43.0 34.6 69.5 66.1 54.7
04.4
08 FoundPose Featuremetric + MegaPose 55.7 51.0 63.3 43.3 35.7 69.7 66.1 55.0
06.4
9 MegaPose [41] MegaPose
49.9
47.7 65.3
36.7
31.5
65.4
60.1
50.9
31.7
With pose refinement (multiple hypotheses):
10 FoundPose* Featuremetric, 5 hyp.
42.0
43.6
60.2
30.5
27.3
53.7
51.3
44.1
07.4
11 FoundPose* MegaPose, 5 hyp.
58.6
54.9
65.7
44.4
36.1
70.3
67.3
56.8
11.2
12 FoundPose* Featuremetric + MegaPose, 5 hyp. 61.0 57.0 69.4 47.9 40.7
72.3
69.0 59.6
20.5
13 GenFlow [57] GenFlow, 5 hyp.
56.3
52.3
68.4
45.3
39.5
73.9
63.3
57.1
20.9
14 GenFlow [57] GenFlow, 16 hyp. 57.2 52.8 68.8 45.8 39.8 74.6
64.2
57.6
40.5
15 MegaPose [41] MegaPose, 5 hyp.
56.0
50.7
68.4
41.4
33.8
70.4
62.1
54.7
47.4
16 MegaPose [41] MegaPose, 10 hyp.
56.0
50.8
68.7
41.9
34.6
70.6
62.0
54.9
53.9
[…]
Compared methods. FoundPose is compared against RGB methods evaluated on the unseen object pose estimation task of the BOP Challenge 2023 [34], GenFlow[57] and MegaPose [41], along with OSOP [82] and ZS6D [3]. FoundPose, GenFlow, MegaPose and ZS6D use the same segmentation masks that were produced by CNOS [60] and provided to the challenge participants, while OSOP relies on a custom detector of unseen objects for LM-O, HB, and YCB-V, and on Mask R-CNN [25] trained for the specific objects for T-LESS (hence we do not include the T-LESS result). MegaPose consists of coarse pose estimation followed by a render-and-compare refiner. Besides variants of FoundPose where the coarse poses are refined by the featuremetric refinement (Sec. 3.5), we evaluate variants where the coarse poses are refined by 5 iterations of the MegaPose refiner.
[…]
4.2 Main results
Accuracy. Among methods that do not apply any refinement stage, FoundPose (without the featuremetric refinement) produces significantly more accurate poses than the competitors, achieving +14 and +17 AR on the seven BOP datasets on average w.r.t. the coarse versions of GenFlow [57] or MegaPose [41] (rows 1–3 in Tab. 1). The featuremetric refinement brings an extra improvement of +5 AR on average (rows 1 vs 6). At an additional computational cost, a large improvement of +17 AR (rows 1 vs 7) can be achieved if the coarse poses from FoundPose are refined by the iterative render-and-compare approach from MegaPose, which is trained on 2M+ synthetic images of diverse objects and proven remarkably effective. When initiated with coarse poses from FoundPose, the MegaPose refiner achieves +4 higher AR score than when initiated with poses from the original coarse pose estimation stage of MegaPose (rows 7 vs 9). Additional improvements at additional computational cost can be achieved if the refinement is applied to multiple pose hypotheses and the top refined pose is reported as the final estimate (rows 10–16). We achieve the overall best average AR score of 59.6 AR when top 5 pose hypotheses are optimized with the featuremetric refinement followed by the MegaPose refinement. On both the single and multi hypotheses setups, the combination of the two refinement approaches achieves the best scores (rows 8 and 12), suggesting their complementarity. This entry outperforms multi-hypotheses versions of MegaPose and GenFlow, including the winning RGB method from BOP Challenge 2023 [34] (row 14).
@@ 4.2 Main results
Backbones for extraction of image patch descriptors:
01 DINOv2 ViT-L – layer 18
39.6
33.8
46.7
23.9
20.4
50.8
45.2
37.2
1.7
2 DINOv2 ViT-L – layer 23
23.2
22.8
31.2
10.3
09.7
33.0
34.0
23.5
1.5
3 DINOv2 ViT-S – layer 9
34.0
31.6
42.7
21.7
16.8
46.8
44.7
34.0
1.3
4 DINOv2 ViT-S – layer 11
22.8
24.2
29.8
11.9
10.5
30.4
36.4
23.7
1.3
5 SAM ViT-L [40] – layer 23
02.2
12.8
09.2
07.5
06.0
10.6
26.9
10.7
3.4
6 DenseSIFT – step size 7px
03.2
02.6
06.5
10.5
02.9
05.6
22.2
07.6
1.4
7 S2DNet[21]
00.8
01.2
00.8
01.4
01.2
01.2
01.3
01.1
1.8
@@ 7 S2DNet[21]
8 Retrieval by cls token
19.9
17.8
24.6
10.3
13.6
17.7
23.6
18.2
1.6
9 Retrieval by cls token with black bg.
25.5
26.2
30.3
16.7
13.6
29.3
34.4
25.1
1.6
@@ 8 Retrieval by cls token
10 Pose given by the top matched template
20.3
18.5
23.0
12.8
12.4
19.6
17.6
17.7
1.0
11 Ground-truth instead of CNOS masks
45.6
53.1
57.1
30.6
50.9
@@ 11 Ground-truth instead of CNOS masks
Table 2: Ablation experiments. We analyze different backbones for extraction of image patch descriptors (rows 1–7), compare our bag-of-words template matching with alternative approaches based on matching the cls token (rows 1 vs 8 and 9), evaluate our template matching approach by taking the pose associated with the top template as the final pose estimate (row 10), and show the accuracy of coarse poses from FoundPose when ground-truth masks are used instead of the CNOS masks[60] (row 11; ground-truth annotations for ITODD and HB are not publicly available).

### v2

@@ 3.2 Template-based object representation
Template generation. Given a texture-mapped 3D object model, we render $n$ RGB-D templates showing the model under different orientations. The orientations are sampled to uniformly cover the $SO(3)$ group of 3D rotations [1], and the model is rendered using a standard rasterization technique [74] with a black background and fixed lighting. The size of templates is $S\times S$ pixels, and the objects are rendered such that the longer side of their 2D bounding box is $\delta S$ pixels long, with $\delta<1$. At inference, we generate crops of the query image with the same size and padding (to allow for errors of segmentation masks around which we crop the image).
@@
Patch descriptors registered in 3D. For each RGB-D template with an index $t\in\{1,\dots,n\}$, we split the RGB channels into $m$ non-overlapping patches of $14\times 14$ pixels and calculate their patch descriptors $\{\mathbf{p}_{t,i}\}_{i=1}^{m}$. A patch descriptor is calculated as $\mathbf{p}_{t,i}=\phi_{d}(\mathbf{p}^{\prime}_{t,i})$, where $\mathbf{p}^{\prime}_{t,i}$ is the raw patch descriptor extracted by DINOv2 and $\phi_{d}:\mathbb{R}^{r}\mapsto\mathbb{R}^{d}$ projects the $r$-dimensional raw descriptor to the top $d$ PCA components, which are calculated from valid patch descriptors of all $n$ templates. A patch is considered valid if its 2D center falls inside the object mask and the PCA-based dimensionality reduction is applied to increase efficiency. Then we represent a template $t$ by a set $T_{t}=\{(\mathbf{p}_{t,j},\mathbf{x}_{j})\;|\;j\in M\}$, where $M$ are indices of valid patches, $\mathbf{x}_{j}$ is a 3D location (in the coordinate space of the 3D object model) whose 2D projection is at the center of patch $j$. The 3D locations, which are calculated from the depth channel of the template and known camera intrinsics, enable establishing 2D-3D correspondences at inference.
@@
Bag-of-words descriptors. At onboarding, we also pre-calculate bag-of-words descriptors of all templates to enable efficient template retrieval at inference using the classical bag-of-words image retrieval technique [76, 66], which mimics text-retrieval systems with the analogy of visual words. Specifically, we define visual words as the centroids of $k$-means clusters of patch descriptors extracted from all templates of an object. To calculate the bag-of-words descriptor of a template $t$, we assign patch descriptors from the template representation $T_{t}$ to the nearest visual words and describe the template by a vector $\mathbf{b}_{t}=(b_{1},b_{2},\dots,b_{k})$. This vector consists of weighted word frequencies defined as $b_{i}=(n_{i,t}/n_{t})\log(N/n_{i})$, where $n_{i,t}$ is the number of occurrences of word $i$ in template $t$, $n_{t}$ is the total number of words in template $t$, and $n_{i}$ is the number of occurrences of word $i$ in all $N$ templates. The first term ($n_{i,t}/n_{t}$) weights words that occur often in a particular template and therefore describe the template well, while the second term ($\log(N/n_{i})$) downweights words that occur often in any template. As visual words generated by clustering may suffer from quantization errors, we follow [66, 86] and soft-assign each patch descriptor to several nearest words with weights defined by $\exp(-d^{2}/2\sigma^{2})$, where $d$ is the Euclidean distance of the descriptor from the word and $\sigma$ is a parameter of the method.
@@ 3.3 Template retrieval by bag-of-words matching
Perspective cropping. At inference, we start by cropping the image region around a given object segmentation mask. To minimize perspective distortion and achieve a crop that resembles a template, we generate the crop by warping the query image to a virtual pinhole camera focused on the segmentation mask. The virtual camera is constructed such that its optical axis passes through the center of the 2D bounding box of the mask, the viewport size is $S\times S$ pixels, and the longer side of the warped 2D bounding box is $\delta S$ pixels long.
Retrieving similar templates. To retrieve a small set of templates, we calculate the bag-of-words descriptor of the crop (as in Sec. 3.2) and calculate its cosine similarity (i.e., normalized scalar product) with bag-of-words descriptors of all object templates. We select $h$ templates with the highest cosine similarity, which provide approximate hypotheses on the object orientation for the subsequent pose estimation stage.
This retrieval technique is efficient and robust to partial occlusions. When an object is partially occluded, its visible part still contributes visual words describing the object. The cosine similarity then normalizes the magnitude and focuses on the direction of the bag-of-words descriptors and is, therefore, less sensitive to the number and more to the type of present words. This robustness has been described in prior work [65, 76] and also in our experiments (see results on LM-O and T-LESS in Sec. 4).
Since the bag-of-words descriptor represents an image as a bag of unordered visual words, the retrieval can be typically improved by re-ranking the results with a spatial verification stage [65]. However, in our case, a similar verification is implicitly done by the subsequent PnP-RANSAC (spatially consistent correspondences are expected to yield a better pose estimate), and bags of words are not actually unordered as the used patch descriptors from an intermediate DINOv2 layer, from which the words are constructed, contains 2D positional information.
Figure 3: Visualization of DINOv2 patch descriptors. Shown are top three PCA components of patch descriptors from different layers of DINOv2 ViT-L[15], for a textured object from YCB-V[91] (top) and a symmetric and texture-less object from T-LESS [31] (bottom). As observed in[2] and also clearly visible in these visuals, the patch descriptors contain gradually less positional and more semantic information when going from shallower to deeper layers – the different coloring of object sides (red left vs yellow right) in Layer 13 gradually blends to a solid color (orange) in Layer 23. FoundPose performs the best with descriptors from layer 18, which presumably provides the right information mix. We observed that these descriptors produce geometrically consistent correspondences even on symmetric and texture-less objects – when the semantic information is ambiguous (due to symmetries or a lack of texture), the positional information prioritizes matching patches from the same object side.
@@ 3.4 Pose estimation from 2D-3D correspondences
Crop-to-template patch matching. For each retrieved template $t$, we match patch descriptors from the crop to the nearest descriptors from $T_{t}$ (in terms of the Euclidean distance), and establish 2D-3D correspondences $C_{t}=\left\{\left(\mathbf{u}_{i},\mathbf{x}_{i}\right)\right\}_{i=1}^{m}$, where $\mathbf{u}_{i}$ is the 2D center of a query patch and $\mathbf{x}_{i}$ is the 3D location associated with the matched patch descriptor from $T_{t}$. The cyclic matching from Goodwin et al. [24] did not help in our setup.
Establishing 2D-3D correspondences by crop-to-template patch matching is a considerably simpler problem than exhaustive matching against patch descriptors from all templates, which would be necessary without the template retrieval stage. Moreover, we demonstrate that the template-based approach can effectively handle arbitrary objects, including challenging objects with symmetries and without a significant texture. The ambiguity of 2D-3D correspondences, for which such objects are notoriously known [30], is eliminated by (1) restricting the candidate patches to only a single template and (2) using patch descriptors from an intermediate layer of DINOv2 which contain both semantic and 2D positional information [2]. We find that the positional information is crucial for producing geometrically consistent correspondences when the semantic information is not discriminative due to symmetries or a lack of texture (see Fig. 3 and Sec. 4.3).
Pose fitting. An object pose $(\mathbf{R}_{t},\mathbf{t}_{t})$, defined by a 3D rotation $\mathbf{R}_{t}$ and a 3D translation $\mathbf{t}_{t}$ from the model space to the camera space, is estimated for each retrieved template $t$ from 2D-3D correspondences $C_{t}$ by solving the Perspective-n-Point (PnP) problem. We solve this problem by the EPnP algorithm [43] combined with the RANSAC fitting scheme [22] for robustness. In this scheme, PnP is solved repeatedly on a randomly sampled minimal set of 4 correspondences, and the final output is defined by the pose hypothesis with the highest quality, which we define by the number of inlier correspondences [22]. From the set of $h$ poses estimated from the $h$ retrieved templates, the pose with the highest quality is selected as the final coarse pose estimate $(\mathbf{R}_{c},\mathbf{t}_{c})$.
@@ 3.5 Featuremetric pose refinement
Reducing 2D-3D discrepancy. The 2D-3D correspondences are established by linking the 2D centers of matched crop and template patches and lifting the centers of the template patches to 3D using the rendered depth channel. Since the patches are relatively large ($14\times 14$ px), the 2D centers of the crop patches may not precisely align with 2D projections of the corresponding 3D points, even when projected using the ground-truth model-to-crop transformation. To compensate for the potential discrepancy, we refine the pose estimates by aligning template patches to their optimal locations in the image crop by a featuremetric optimization described below.
Featuremetric alignment. The coarse pose $(\mathbf{R}_{c},\mathbf{t}_{c})$ is refined by Levenberg-Marquardt (L-M) [44, 52], an iterative non-linear optimization algorithm. Each L-M iteration updates the pose parameters by a gradient-descent step, minimizing a real-valued cost while adaptively selecting between the first and second-order gradients depending on the cost value. We initialize the optimization with the coarse pose $(\mathbf{R}_{c},\mathbf{t}_{c})$ and obtain a refined pose $(\mathbf{R}_{r},\mathbf{t}_{r})$ by minimizing the following featuremetric error:
@@
where $\rho$ is a robust cost function by Barron [6], $(\mathbf{p}_{i},\mathbf{x}_{i})\in T_{t}$ is a descriptor and the corresponding 3D location of patch $i$ from template $t$, and $\pi:\mathbb{R}^{3}\mapsto\mathbb{R}^{2}$ represents the 2D projection. The feature map $\textbf{F}_{q}\in\mathbb{R}^{a\times a\times d}$ is obtained by stacking the patch descriptors of the query image and is sampled with bilinear interpolation at normalized 2D projections. The spatial resolution of the feature map is $a\times a$ with $a=S/s$, where $s\times s$ is the patch size. Note that a similar featuremetric alignment was applied in several works [77, 73], typically in combination with features trained specifically for L-M. In our work, we apply the alignment directly on DINOv2 features without any training. Even though there is no guarantee for reaching the global optimum, as in other methods optimizing photo/featuremetric objectives, we observed that a plausible solution is often reached.
@@ 4 Experiments
Figure 4: Example FoundPose results on HB, LM-O, IC-BIN, TUD-L, ITODD and T-LESS datasets, showing that our method can handle a broad range or objects, including textured, texture-less and symmetric ones. Each example shows the query image crop with the CNOS mask in white (top left), retrieved templates (middle row), matched patch descriptors of the crop and the template that led to the top-quality pose estimate (bottom row), and the contour of the ground-truth pose in red, the coarse pose in blue, and the refined pose in green (top right).
@@ 4.1 Experimental setup
Compared methods. FoundPose is compared against model-based RGB methods evaluated on the unseen object pose estimation task of the BOP Challenge 2023 [34]: GenFlow[54], MegaPose [41], GigaPose [59], and also against ZS6D [3] and OSOP [75]. Except for OSOP, all of these methods (including FoundPose) use the same segmentation masks that were produced by CNOS [58] and provided to the challenge participants. OSOP relies on a custom detector of unseen objects for LM-O, HB, and YCB-V, and on Mask R-CNN [25] trained for specific objects for T-LESS (hence we do not include the T-LESS result). Besides variants of FoundPose where the coarse poses are refined by the featuremetric refinement (Sec. 3.5), we evaluate variants with poses refined by 5 iterations of the MegaPose refiner (i.e., the last stage of the MegaPose pipeline [41]).
[…]
4.2 Main results
Accuracy. Among methods that do not apply any refinement stage, FoundPose (without the featuremetric refinement) produces significantly more accurate poses than the competitors, achieving +10, +14, and +16 AR on the seven BOP datasets on average compared to the coarse versions of GigaPose [59], GenFlow [54], and MegaPose [41] (rows 1–4 in Tab. 1). The featuremetric refinement brings an extra improvement of +5 AR on average (rows 1 vs 7). At an additional computational cost, a large improvement of +17 AR (rows 1 vs 8) can be achieved if the coarse poses from FoundPose are refined by the iterative render-and-compare approach from MegaPose, which is trained on 2M+ synthetic images of diverse objects and proven remarkably effective. When initiated with coarse poses from FoundPose, the MegaPose refiner achieves +4 higher AR score than when initiated with poses from the original coarse pose estimation stage of MegaPose (rows 8 vs 11). Further improvements at further computational cost can be achieved if the refinement is applied to multiple pose hypotheses and the top refined pose is reported as the final estimate (rows 12–17). We achieve the overall best average AR score of 59.6 AR when top 5 pose hypotheses (generated from 5 retrieved templates) are optimized with the featuremetric refinement followed by the MegaPose refinement. On both the single and the multi hypotheses setups, combining the two refinement approaches achieves the best scores (rows 9 and 14), suggesting their complementarity. This entry outperforms multi-hypotheses versions of MegaPose and GenFlow (rows 16 and 17), which are the top-performing RGB methods from the BOP Challenge 2023 [34], as well as GigaPose[59] with the MegaPose refinement (row 15).
# Method Pose refinement
No train.
@@ 4.2 Main results
01 FoundPose 39.6 33.8 46.7 23.9 20.4 50.8 45.2 37.2 01.7
2 GigaPose [59] 29.9 27.3 30.2 23.1 18.8 34.8 29.0 27.6 00.9
3 GenFlow [54] 25.0 21.5 30.0 16.8 15.4 28.3 27.7 23.5 03.8
4 MegaPose [41] 22.9 17.7 25.8 15.2 10.8 25.1 28.1 20.8 15.5
5 OSOP [75] 31.2 49.2 33.2
6 ZS6D [3] 29.8 21.0 32.4
@@
07 FoundPose Featuremetric 39.5 39.6 56.7 28.3 26.2 58.5 49.7 42.6 02.6
08 FoundPose MegaPose 55.4 51.0 63.3 43.0 34.6 69.5 66.1 54.7 04.4
09 FoundPose Feat. + MegaPose 55.7 51.0 63.3 43.3 35.7 69.7 66.1 55.0 06.4
10 Gigapose [59] MegaPose 55.6 54.6 57.8 44.3 37.8 69.3 63.4 54.7 02.4
11 MegaPose [41] MegaPose 49.9 47.7 65.3 36.7 31.5 65.4 60.1 50.9 31.7
With pose refinement (5 hypotheses):
12 FoundPose Featuremetric 42.0 43.6 60.2 30.5 27.3 53.7 51.3 44.1 07.4
13 FoundPose MegaPose 58.6 54.9 65.7 44.4 36.1 70.3 67.3 56.8 11.2
14 FoundPose Feat. + MegaPose 61.0 57.0 69.4 47.9 40.7 72.3 69.0 59.6 20.5
15 GigaPose [59] MegaPose 59.9 57.0 64.5 46.7 39.7 72.2 66.3 57.9 07.3
16 GenFlow [54] GenFlow 56.3 52.3 68.4 45.3 39.5 73.9 63.3 57.1 20.9
17 MegaPose [41] MegaPose 56.0 50.7 68.4 41.4 33.8 70.4 62.1 54.7 47.4
Table 1: Performance on the seven core BOP datasets [82]. The table shows Average Recall (AR) scores per dataset, the average AR score, and the time to estimate poses of all objects in an image averaged over the datasets (in seconds). Methods for coarse pose estimation (without applying any refinement) are in the upper part, methods applying a refinement stage on a single pose hypothesis are in the middle, and methods refining multiple pose hypotheses and reporting the top refined pose (with the highest refinement score) are at the bottom. Methods that do not undergo any task-specific training are marked with .
# Method
@@
Backbones for extracting patch descriptors:
01 DINOv2 ViT-L – layer 18 39.6 33.8 46.7 23.9 20.4 50.8 45.2 37.2 1.7
2 DINOv2 ViT-L – layer 23 23.2 22.8 31.2 10.3 09.7 33.0 34.0 23.5 1.5
3 DINOv2 ViT-S – layer 9 34.0 31.6 42.7 21.7 16.8 46.8 44.7 34.0 1.3
4 DINOv2 ViT-S – layer 11 22.8 24.2 29.8 11.9 10.5 30.4 36.4 23.7 1.3
5 SAM ViT-L [40] – layer 23 02.2 12.8 09.2 07.5 06.0 10.6 26.9 10.7 3.4
6 DenseSIFT – step size 7px 03.2 02.6 06.5 10.5 02.9 05.6 22.2 07.6 1.4
7 S2DNet[23] 00.8 01.2 00.8 01.4 01.2 01.2 01.3 01.1 1.8
@@
8 Retrieval by cls token 19.9 17.8 24.6 10.3 13.6 17.7 23.6 18.2 1.6
9 Retrieval by cls token with black bg. 25.5 26.2 30.3 16.7 13.6 29.3 34.4 25.1 1.6
@@
10 Pose given by the top matched template 20.3 18.5 23.0 12.8 12.4 19.6 17.6 17.7 1.0
11 Ground-truth instead of CNOS masks 45.6 53.1 57.1 30.6 50.9
@@
Table 2: Ablation experiments. We analyze different backbones for extracting patch descriptors (rows 1–7), compare our bag-of-words template retrieval with alternative approaches based on matching the cls token (rows 1 vs 8 and 9), evaluate our retrieval approach by considering the pose associated with the top template as the final pose estimate (row 10), and show the accuracy of coarse poses from FoundPose when ground-truth masks are used instead of the CNOS masks[58] (row 11; ground-truth annotations for ITODD and HB are not publicly available).